40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Jun 29 02:17:40 2012, yuta, Update, IOO, prep for the vent - beam attenuating 
    Reply  Fri Jun 29 03:21:32 2012, yuta, Update, General, prep for the vent - others MEDMscreenshotswithCOW_20120629.png
       Reply  Fri Jun 29 11:02:00 2012, steve, Update, VAC, the vent is at 500 torr sussum.pngpzt.png
          Reply  Fri Jun 29 13:11:31 2012, steve, Update, VAC, 40m IFO at atm vent06292012.png40m@ATM.png
Message ID: 6892     Entry time: Fri Jun 29 02:17:40 2012     Reply to this: 6893
Author: yuta 
Type: Update 
Category: IOO 
Subject: prep for the vent - beam attenuating 

[Koji, Jamie, Yuta]

We attenuated the incident beam (1.2 W -> 11 mW) to the vacuum chamber to be ready for the vent.
The beam spot on the MC mirrors didn't changed significantly, which means the incident beam was not shifted so much.

What we did:
 1. Installed HWP, PBS(*) and another HWP between the steering mirrors on PSL table for attenuating the beam. We didn't touched steering mirrors(**), so the incident beam to the IFO should be recovered easily, by just taking HWPs and PBS away. The power to the MC was reduced from 1.2 W to 11 mW.

(*) We stole PBSO from the AS AUX laser setup.
(**) Actually, we accidentally touched one of the steering mirrors, but we recovered them. We did the recovery tweaking the touched nob and minimizing the MC reflection. We confirmed the incident beam was recovered by measuring MC beamspot positions(below).

 2. Aligned PBS by minimizing MC reflection, adjusted first HWP so that the incident beam will be ~10 mW, and adjusted last HWP to minimize MC reflection (make the incident beam to the MC be p-polarization).

 3. To do the alignment and adjusting, we put 100% reflection mirror (instead of 10% BS) for the MC reflection PD to increase the power to the PD. That means, we don't have MC WFS right now.

 4. Tweaked MC servo gains to that we can lock MC in low power mode. It is quite stable right now. We didn't lose lock during beam spot measurement.

 5. Measured beam spot positions on the MC mirrors and convinced that the incident beam was not shifted so much (below). They look like they moved ~0.2 mm, but it is with in the error of the MC beam spot measurement.

# filename      MC1pit  MC2pit  MC3pit  MC1yaw  MC2yaw  MC3yaw  (spot positions in mm)
./dataMCdecenter/MCdecenter201206281154.dat     3.193965        4.247243        2.386126        -6.639432       -0.574460       4.815078    this noon
./dataMCdecenter/MCdecenter201206282245.dat     3.090762        4.140716        2.459465        -6.792872       -0.651146       4.868740    after recovered steering mirrors
./dataMCdecenter/MCdecenter201206290135.dat     2.914584        4.240889        2.149244        -7.117336       -1.494540       4.955329    after beam attenuation

 6. Rewrote matlab code sensemcass.m to python script sensemcass.py. This script is to calculate beam spot positions from the measurement data(see elog #6727). I think we should make senseMCdecenter script better, too, since it takes so much time and can't stop and resume the measurement if MC is unlocked.

ELOG V3.1.3-